1. Homer is a street person whose preferences over the two goods he consumes, food (good 1) and

fortified wine (good 2) are represented by the utility function U(x1,x2)=XjX28 (MUi=x2", and

MU2=8xıx27). Homer receives a pension of $100.00 per week. He can buy food at the local

dispensary at a price of $4 a bucket. The local convenience store sells fortified wine at $6 a gallon

jug.

What is Homer's MRS?


1
Expert's answer
2021-09-15T11:28:45-0400

"Given:\\

U=x_1x2^8\\

MU_1=x2^8\\

MU_2=8x_1x_2^7\\

MRS_{12}=\frac{MU_1}{MU_2}\\

MRS_{12}=\frac{x_2^8}{8x_1x_2^7}\\

MRS_{12}=\frac{x_0^2}{8x_1}\\

So, Homer's MRS is \frac{x_2}{8x_1}"